# DualBEV: CNN is All You Need in View Transformation

### v1

@@ Abstract
Camera-based Bird’s-Eye-View (BEV) perception often struggles between adopting 3D-to-2D or 2D-to-3D view transformation (VT). The 3D-to-2D VT typically employs resource-intensive Transformer to establish robust correspondences between 3D and 2D feature, while the 2D-to-3D VT utilizes the Lift-Splat-Shoot (LSS) pipeline for real-time application, potentially missing distant information. To address these limitations, we propose DualBEV, a unified framework that utilizes a shared CNN-based feature transformation incorporating three probabilistic measurements for both strategies. By considering dual-view correspondences in one-stage, DualBEV effectively bridges the gap between these strategies, harnessing their individual strengths. Our method achieves state-of-the-art performance without Transformer, delivering comparable efficiency to the LSS approach, with 55.2% mAP and 63.4% NDS on the nuScenes test set. Code will be released at https://github.com/PeidongLi/DualBEV.
@@
View Transformation BEV CNN Autonomous Driving
| HeightFormer | FB-BEV | DualBEV |
@@ 1 Introduction
Effective BEV object detection in autonomous driving relies on precise feature transformation from the perspective to the BEV space, facilitated by the VT module. Current methods predominantly employ either a 2D-to-3D or 3D-to-2D strategy. In 2D-to-3D methods[7, 14, 13, 8, 34, 10, 38], dense 2D features are elevated by predicting depth probabilities, but the inherent uncertainty in depth prediction can introduce inaccuracies, particularly in distant regions. Conversely, 3D-to-2D methods[30, 19, 15, 9, 36, 32, 17, 18] often use 3D queries to sample 2D features, leveraging Transformer[29] to learn attention weights for each correspondence between 3D and 2D features, while introducing computational and deployment complexities.
@@
In this paper, we argue that both VT inherently establish correspondences between 3D and 2D features from different views, while LSS and Transformer serving as distinct methods for evaluating these correspondences. To unify the dual VT, we propose a unified feature transformation (illustrated in Fig. 1) applicable to both 2D-to-3D and 3D-to-2D VT, evaluating the correspondences by three probabilistic measurements: 1) BEV Probability, aimed at mitigating the impact of blank BEV grids in feature construction; 2) Projection Probability, which distinguishes multiple correspondences, accounting for different 3D points projecting into the same 2D position; 3) Image Probability, aiding in excluding background features during feature transformation.
@@
Combining HeightTrans and Prob-LSS, our research introduces DualBEV (see Tab. 1), an innovative approach that incorporates and considers correspondences from both BEV and perspective view in the one-stage manner, eliminating the reliance on initial features. Furthermore, we propose a robust BEV feature fusion module termed Dual Feature Fusion (DFF) module. This module enhances the integration of dual BEV features by leveraging channel attention module, while spatial attention module further helps to refine BEV probability prediction. DualBEV operates on the principle of ”broad input, stringent output”, leveraging precise dual-view probabilistic correspondences to comprehend and represent the probability distribution of the scene.
@@ 3.1.2 Prob-Sampling
To distinguish multiple 3D points hitting the same 2D position, we further use a projection probability obtained by the trilinear grid sampler in the depth map to evaluate these multiple correspondences.
@@ 3.1.3 Acceleration
As showed in Tab. 1, this representation bears similarity to BEV Pooling. Therefore, we can establish the lookup table by projecting predefined 3D points into the image space and subsequently calculating the index in the feature map and depth map.Unlike BEV Pooling, the index in BEV map is constant in our method. Then we can use the same CUDA operator as in BEVPoolv2[5] with the lookup table to speed up the calculation of the BEV features during inference.
@@ 3.3 Dual Feature Fusion
Unifying Eq. 7 and Eq. 8, we observe that BEV probability can be obtained after the fusion of dual features. Inspired by the CBAM[31] and AFF[2] framework, we propose a Dual Feature Fusion (DFF) module to integrate these features and predict the BEV probability effectively. The DFF module comprises a fusion module , which utilizes channel attention to predict weights for dual feature summation, and a Spatial Attention Enhanced ProbNet (SAE-ProbNet) to predict the BEV probability , formulated as:
@@ 3.3.1 Fusion Module
As showed in Fig. 4, focuses more on close objects occupying most of the image pixels, while tends to contain information of distant objects. This observation suggests two-stage methods[32, 16] initialized by LSS features, may not guide second-stage VT to extract features at distant regions.
@@ 3.3.2 BEV Probability Prediction
As depicted in Fig. 5, tthe local stream, ProbNet, employs a 33 convolution kernel to reduce the channel dimensions. Subsequently, it undergoes processing through a ResBlock-CBAM[31] and a 11 convolution operation to acquire local attention. Supervised by a BEV mask, ProbNet utilizes BCE loss and Dice loss[24], with BEV Centerness[35] also incorporated into the loss to encourage the network to focus more on distant objects. On the other hand, the global stream utilizes a 77 convolution kernel to compute the average and maximum values of the input feature, thus enlarging the perception field in BEV space. These streams are then combined before passing through the sigmoid function . This design aims to equip our module with the capability to capture both local and global attention for enhanced BEV probability prediction.
@@ 4.3 Main Results
We adopt BEVDepth[14] with BEVPoolv2[5] as baseline method. As illustrated in Tab. 2, DualBEV outperforms the baseline method with +1.0% mAP and +1.8% NDS in single frame. For multi-frame, DualBEV not only surpasses the baseline with +1.2% mAP and +1.9% NDS, but also outperforms BEVStereo[13] which uses temproral-stereo module for precise depth estimation. By leveraging HT features which capture the distant information, and efficient fusion with LSS feature, we also have a 0.5% NDS increase compared to SA-BEV[10] and a 0.6% NDS increase compared to FB-BEV[16].
@@
Tab. 3 demonstrates that DualBEV achieves state-of-the-art performance, yielding a remarkable 55.2% mAP and 63.4% NDS compared to prior works in VT. Notably, our approach exhibits a significant margin of 1.9% in mAP over SA-BEV[10] and 1.5% over FB-BEV[16], while also surpassing them by 1.0% in NDS. Additionally, when compared to SOLOFusion[25], which focus on temporal module, DualBEV maintains superiority with a lead of 1.2% in mAP and 1.5% in NDS even without temproral-stereo module. Furthermore, DualBEV outperforms the previous state-of-the-art Transformer-based VT approach SparseBEV[18], by a notable margin of 0.9% in mAP and 0.7% in NDS, proving the pivatol role of unveiling precise correspondences in VT.
@@ 4.4.1 Impact of Each Component
We first explore each component’s impact in Tab. 4(a). Starting with BEVDepth[14] as the baseline, the initial integration of ProbNet to provide BEV probability results in a 0.4% increase in both mAP and NDS. Combining HT features led to a 0.4% mAP and 0.4% NDS increase, showcasing compensation for Prob-LSS. Subsequently, we employ the CAF module to softly select features from the two streams, resulting in a further 0.3% mAP and 0.1% NDS increase. Finally, by incorporating the SAE module for improved BEV probability prediction with the addition of spatial attention module alongside ProbNet, our method achieves a substantial 0.9% improvement in NDS. The overall components gain a 1.0% mAP and 1.8% NDS improvement. Additionally, we also test the latency of VT from image feature to BEV feature including depth estimation. The entire proposed components reveals an extra latency of merely 1.5ms, with ProbNet introducing 0.9ms of this total.
@@ 4.4.4 Effect of Sampling Strategy in Height
In HeightTrans, we apply a multi-resolution (MR) sampling strategy within the height range. We evaluate different sampling strategies as shown in Tab. 4(d). The performance almost reaches saturation after 8 uniformly sampled points. Notably, our proposed approach achieves a 0.6% increase in NDS compared to 4 uniform points in BEVFormer[15], and even a 0.2% enhancement compared to 16 sampling points.
@@ 4.4.6 Effect of Fusion Strategy
Tab. 4(f) compares different fusion strategy for HT feature and Prob-LSS feature. We first explore the refinement strategy employed in FB-BEV[16], which utilizes Prob-LSS features to predict the BEV probability for HeightTrans. However, the refinement strategy yielded even worse results than directly summing the both features for BEV probability prediction, with a 0.4% NDS decrease. Our proposed DFF strategy achieved a notable improvement, surpassing the refinement strategy by 0.8% in NDS.

### v2

# DualBEV: Unifying Dual View Transformation with Probabilistic Correspondences

@@ Abstract
Camera-based Bird’s-Eye-View (BEV) perception often struggles between adopting 3D-to-2D or 2D-to-3D view transformation (VT). The 3D-to-2D VT typically employs resource-intensive Transformer to establish robust correspondences between 3D and 2D features, while the 2D-to-3D VT utilizes the Lift-Splat-Shoot (LSS) pipeline for real-time application, potentially missing distant information. To address these limitations, we propose DualBEV, a unified framework that utilizes a shared feature transformation incorporating three probabilistic measurements for both strategies. By considering dual-view correspondences in one stage, DualBEV effectively bridges the gap between these strategies, harnessing their individual strengths. Our method achieves state-of-the-art performance without Transformer, delivering comparable efficiency to the LSS approach, with 55.2% mAP and 63.4% NDS on the nuScenes test set. Code is available at https://github.com/PeidongLi/DualBEV.
@@
View Transformation BEV Autonomous Driving
@@ 1 Introduction
Effective BEV object detection in autonomous driving relies on precise feature transformation from the perspective to the BEV space, facilitated by the VT module. Current methods predominantly employ either a 2D-to-3D or 3D-to-2D strategy. In 2D-to-3D methods [7, 14, 13, 8, 34, 10, 38], dense 2D features are elevated by predicting depth probabilities, but the inherent uncertainty in depth prediction can introduce inaccuracies, particularly in distant regions. Conversely, 3D-to-2D methods [30, 19, 15, 9, 36, 32, 17, 18] often use BEV queries to sample 2D features, leveraging Transformer [29] to learn attention weights for each 3D-2D correspondence, while introducing computational and deployment complexities.
@@
In this paper, we argue that both VT inherently establish correspondences between 3D and 2D features from different views, while LSS and Transformer serving as distinct methods for evaluating these correspondences. To unify dual VT, we propose a unified feature transformation (illustrated in Fig. 1) applicable to both 2D-to-3D and 3D-to-2D VT, evaluating the correspondences by three probabilistic measurements: 1) BEV Probability, aimed at mitigating the impact of blank BEV grids in feature construction; 2) Projection Probability, which distinguishes multiple correspondences, accounting for different 3D points projecting into the same 2D position; 3) Image Probability, aiding in excluding background features during feature transformation.
@@
Combining HeightTrans and Prob-LSS, our research introduces DualBEV (see Tab. 1), an innovative approach that incorporates and considers correspondences from both BEV and perspective view in the one-stage manner, eliminating the reliance on initial features. Furthermore, we propose a robust BEV feature fusion module termed Dual Feature Fusion (DFF) module. This module enhances the integration of dual BEV features by leveraging channel attention module, while spatial attention module further helps to refine BEV probability prediction. DualBEV operates on the principle of "broad input, stringent output", leveraging precise dual-view probabilistic correspondences to comprehend and represent the probability distribution of the scene.
@@ 3.1.2 Prob-Sampling
To distinguish multiple 3D points hitting the same 2D position, we further use a projection probability to evaluate these multiple correspondences. is obtained by a trilinear grid sampler on the depth map , treating the depth channel as the third dimension.
@@ 3.1.3 Acceleration
As shown in Tab. 1, this representation bears similarity to BEV Pooling. Therefore, we can establish the lookup table by projecting pre-defined 3D points into the image space and subsequently calculating the index in the feature map and depth map. Unlike BEV Pooling, the index in BEV map is constant in our method. Then we can use the same CUDA operator as in BEVPoolv2[5] with the lookup table to speed up the calculation of BEV features during inference.
@@ 3.3 Dual Feature Fusion
Unifying Eqs. 7 and 8, we observe that BEV probability can be obtained after the fusion of dual features. Inspired by the CBAM[31] and AFF[2] framework, we propose a Dual Feature Fusion (DFF) module to integrate these features and predict the BEV probability effectively. The DFF module comprises a fusion module , which utilizes channel attention to predict weights for dual feature summation, and a Spatial Attention Enhanced ProbNet (SAE-ProbNet) to predict the BEV probability , formulated as:
@@ 3.3.1 Fusion Module
As shown in Fig. 4, focuses more on close objects occupying most of the image pixels, while tends to contain information of distant objects. This observation suggests two-stage methods[32, 16] initialized by LSS features, may not guide second-stage VT to extract features at distant regions.
@@ 3.3.2 BEV Probability Prediction
As depicted in Fig. 5, the local stream, ProbNet, employs a 33 convolution kernel to reduce the channel dimensions. Subsequently, it undergoes processing through a ResBlock-CBAM[31] and a 11 convolution operation to acquire local attention. Supervised by a BEV mask, ProbNet utilizes BCE loss and Dice loss[24], with BEV Centerness[35] also incorporated into the loss to encourage the network to focus more on distant objects. On the other hand, the global stream utilizes a 77 convolution kernel to compute the average and maximum values of the input feature, thus enlarging the perception field in BEV space. These streams are then combined before passing through the sigmoid function . This design aims to equip our module with the capability to capture both local and global attention for enhanced BEV probability prediction.
@@ 4.3 Main Results
We adopt BEVDepth[14] with BEVPoolv2[5] as baseline method. As illustrated in Tab. 2, DualBEV outperforms the baseline method with +1.0% mAP and +1.8% NDS in single frame. For multi-frame, DualBEV not only surpasses the baseline with +1.2% mAP and +1.9% NDS, but also outperforms BEVStereo[13] which uses temporal-stereo module for precise depth estimation. By leveraging HT features which capture the distant information, and efficient fusion with LSS feature, we also have a 0.5% NDS increase compared to SA-BEV[10] and a 0.6% NDS increase compared to FB-BEV[16].
@@
Tab. 3 demonstrates that DualBEV achieves state-of-the-art performance, yielding a remarkable 55.2% mAP and 63.4% NDS compared to prior works in VT. Notably, our approach exhibits a significant margin of 1.9% in mAP over SA-BEV[10] and 1.5% over FB-BEV[16], while also surpassing them by 1.0% in NDS. Additionally, when compared to SOLOFusion[25], which focus on temporal module, DualBEV maintains superiority with a lead of 1.2% in mAP and 1.5% in NDS even without temporal-stereo module. Furthermore, DualBEV outperforms the previous state-of-the-art Transformer-based VT approach SparseBEV[18], by a notable margin of 0.9% in mAP and 0.7% in NDS, proving the pivatol role of unveiling precise correspondences in VT.
@@ 4.4.1 Impact of Each Component
We first explore each component’s impact in Tab. 4(a). Starting with BEVDepth[14] as the baseline, the initial integration of ProbNet to provide BEV probability results in a 0.4% increase in both mAP and NDS. Combining HT features led to a 0.4% mAP and 0.4% NDS increase, showcasing compensation for Prob-LSS. Subsequently, we employ the CAF module to softly select features from the two streams, resulting in a further 0.3% mAP and 0.1% NDS increase. Finally, by incorporating the SAE module for improved BEV probability prediction with the addition of spatial attention module alongside ProbNet, our method achieves a substantial 0.9% improvement in NDS. The overall components gain a 1.0% mAP and 1.8% NDS improvement. Additionally, we also test the latency of VT from image feature to BEV feature including depth estimation. The entire proposed components reveal an extra latency of merely 1.5ms, with ProbNet introducing 0.9ms of this total.
@@ 4.4.4 Effect of Sampling Strategy in Height
In HeightTrans, we apply a multi-resolution (MR) sampling strategy within the height range. We evaluate different sampling strategies as shown in Tab. 4(d). The performance almost reaches saturation after 8 uniform points. Notably, our proposed approach achieves a 0.6% increase in NDS compared to 4 uniform points utilized in BEVFormer[15], and even a 0.2% enhancement compared to 16 uniform points along the height.
@@ 4.4.6 Effect of Fusion Strategy
Tab. 4(f) compares different fusion strategy for HT feature and Prob-LSS feature. We first explore the refinement strategy employed in FB-BEV[16], which utilizes Prob-LSS features to predict the BEV probability for HeightTrans. However, the refinement strategy yields even worse results than directly summing the both features for BEV probability prediction, with a 0.4% NDS decrease. Our proposed DFF strategy achieves a notable improvement, surpassing the refinement strategy by 0.8% in NDS.
@@ 5 Conclusion and Limitation
Acknowledgements
The authors would like to thank reviewers for their detailed comments and instructive suggestions. This research is supported by National Key R&D Program of China (2022YFB4300300).